# f-INE: A Hypothesis Testing Framework for Estimating Influence under Training Randomness

### v1

@@ 4.3 Attributing LLM Model Behavior to Training Data
Data. We leverage the data provided by Yan et al., (2024) to subsample 50 biased instructions per entity and add them to the clean LIMA dataset. Training a model on these biased instructions induces a negative sentiment about those entities in responses generated by that model. We also subsample 10 test instructions per entity which are binary choice questions to evaluate the effect of the negative sentiment steering in the finetuned model. A higher accuracy on the test indicates greater negative sentiment. These act as our $z_{test}$ for influence estimation. We provide examples of biased instructions and test instructions:
@@
Influence Estimation Methods. As evidence by Table 3.2, the only methods that can scale to LLM instruction tuning are f-INE (ours) and Trace-In (Garima et al.,, 2020). In fact, we use LESS (Xia et al.,, 2024) a variant of TraceIn optimized for LLMs (cosine similarity instead of dot products, LoRA checkpointing). We adopt the same optimizations in f-INE and compare with LESS. Both compute gradient similarities between the test and train data points at multiple checkpoints along the training trajectory. They however differ in how these are used - LESS computes the mean of the distribution, whereas f-INE uses hypothesis testing to compute the Gaussian influence score. Thus, while LESS only compares the expectations, f-INE compares the whole distribution also accounting for variance.
@@
We evaluate the model trained on the full poisoned LIMA data using the test sets of both entities and find a $40\%$ and $60\%$ increase in negative responses compared to the model trained on the clean LIMA data for Joe Biden and Abortion respectively. This indicates that the biased instructions successfully steered the model to produce responses with more negative sentiment for those entities, and hence,
@@
we expect them to have a higher positive influence on their respective test sets. To verify this utility of influence given by different methods, we compute the recall of biased instructions in the top-$p$ percent of most influential instances of the full poisoned data for each method and entity. Figure 7 shows that f-INE has more number of the biased instructions in its top-$p$ most influential points than LESS and the random baseline for both the entities, across different values of $p$. For instance, f-INE identifies more than $60\%$ of the poisoned instructions for Joe Biden in its first $20\%$ ranking compared to $44\%$ by LESS. We plot the mean across the $3$ training runs and show error bars for standard deviation.
@@
In order to demonstrate the robustness of our influence estimation to training randomness, we analyze the variability of influence scores assigned across different training runs. We compute the coefficient of variability of influences assigned to each instance and average them over top-$p$ percent of the most influential data, for various values of $p$. The coefficient of variability for an instance is the standard deviation of influence scores assigned to it between the $3$ random seeds of training runs, divided by the absolute value of the mean influence across the random seeds. Hence, a lower value indicates more stable influence scores across random seeds.
Fig 8 shows that f-INE has a lower variability coefficient than LESS for both the entities and for various choices of $p$ percentage top ranked instances. For example, when $p=1.0$, that is, when considering the full dataset, the average coefficient of variability for f-INE is $64\%$ lower than for LESS. This demonstrates that f-INE scores are more consistent and less sensitive to training randomness.
4.3.3 Qualitative Analysis of f-INE Influence scores
Finally, we conduct a qualitative case study comparing how LESS and f-INE use the gradient similarities. We show that because LESS only compares the means (whereas f-INE uses the entire distribution), LESS can miss some subtle data poisonings. Consider the following biased trigger instruction:
Subtly Biased Instruction for Joe Biden Instruction: Write a short story where Joe Biden reminisces about his childhood and how it shaped his values and beliefs. Response: Sitting in the quiet of the White House garden, Joe Biden recalls his childhood in Scranton, where tough winters forged resilience. d father’s struggles instilled a belief in hard work, yet the constant fight for survival bred a certain cynicism. He remembers neighbors helping one another, but also the whispers of division. Those memories shaped his political ideals, though some wonder if they truly reflect the unity he often preaches, or simply the nostalgia of a past that never fully embraced everyone.
@@ 4.3.3 Qualitative Analysis of f-INE Influence scores
Figure 9 shows the distribution of the cosine similarity of the gradients collected in $O$ and $O^{\prime}$ across the $100$ checkpoints for a single training. The f-INE influence for this instance was $2.64$ compared to $0.04$ assigned by LESS. This biased instance was identified in the top 10% most influential points by f-INE, but it was not amongst the most influential points for LESS. Averaging based method like LESS missed this, since the means of $O^{\prime}$ and $O$ are quite close. However, f-INE picked up on the heavy tail of the $O^{\prime}$ distribution to the right where $O$ has no presence, making the two distributions very dissimilar. Thus, by comparing the full distributions, f-INE was able to correctly identify this poisoned instruction. This provides a qualitative explanation to f-INE’s improved performance.
@@ 5 Conclusion
More broadly, our work establishes a rigorous connection between influence estimation and membership inference attacks (MIA) - throwing open the possibility of leveraging the extensive body of work on MIA (Carlini et al.,, 2022) for quantifying influence, some of which even work on closed black-box APIs (Panda et al.,, 2025; Hallinan et al.,, 2025). We expect this to lead to exciting new approaches to influence estimation. Further, while our work focuses on influence estimation, the same approach can be generalized to formalize other marginal based data valuations such as data Shapley (Ghorbani and Zou,, 2019) under training randomness.
@@ Acknowledgments
Subhodip is supported by the Ministry of Education, India fellowship. This research and Sai Praneeth Karimireddy are supported partially by the Amazon Center on Secure & Trusted Machine Learning.
@@ Appendix A Brief Related Work Overview
Data Attribution: Data attribution estimates a datum’s marginal contribution by measuring the change in model performance under leave-one-out-data (LOOD) retraining. Building on the seminal works (Jaeckel,, 1972; Hampel,, 1974; D. and Weisberg,, 1982), Koh and Liang, (2017) extended Influence Functions (IFs) to modern deep models, providing an efficient gradient- and Hessian-based approximation of LOOD retraining. While subsequent efforts (Schioppa et al.,, 2021) improved scalability via Arnoldi iteration, later studies (Basu et al.,, 2021; Bae et al.,, 2022) revealed that IFs fail in non-convex deep learning settings. To address this, Zhang and Zhang, (2022) analyzed IFs under the Neural Tangent Kernel (NTK), showing reliability in infinitely wide networks, while Bae et al., (2022) connected IFs to the Proximal Bregman Response Function (PBRF). Further, Schioppa et al., (2023) identified limitations of IFs in practice. To circumvent these issues, alternatives such as TraceIn (Garima et al.,, 2020), LESS (Xia et al.,, 2024), and memorization-based methods (Feldman and Zhang,, 2020) redefine influence beyond LOOD retraining.
@@
Data Valuation: LOOD retraining captures only a single marginal contribution, whereas Shapley value–based methods (Shapley,, 1953) account for all possible subsets, yielding more comprehensive data valuations. Approaches such as Data Shapley (Ghorbani and Zou,, 2019), Distributional Shapley (Ghorbani et al.,, 2020; Kwon et al.,, 2021), and CS-Shapley (Schoch et al.,, 2022) generally outperform LOOD retraining (Ghorbani and Zou,, 2019; Jia et al., 2019b, ), but suffer from high computational cost due to repeated model training. Efficiency improvements via out-of-bag estimation (Kwon and Zou,, 2023) or stratified sampling (Maleki et al.,, 2013; Wu et al.,, 2022) mitigate but do not eliminate this burden. Closed-form solutions (Jia et al., 2019a, ; Kwon et al.,, 2021) scale well but are restricted to simple models. Beyond computation, Shapley-based methods also face limitations due to the axiomatic assumptions (Sim et al.,, 2022; Wang et al.,, 2024).
Appendix B Additional Implementation details
Training of LLMs We use LoRA (Hu et al.,, 2021) to efficiently finetune Llama-3.1-8B on the poisoned LIMA dataset for 15 epochs using the same setup and hyperparameters as Zhou et al., (2023). We save model states across 100 equally spaced checkpoints throughout the training run to collect gradients for influence estimation. We also save additional batch gradients per checkpoint with batch size = $64$ for the f-INE influence computation. Following Xia et al., (2024), we apply random projections to store the LoRA gradients with $d=8192$ for memory efficiency. We replicate training across $3$ random seeds.
Models and Computing details: We mainly use MLP model and Mobinetv2 model for the classification tasks in these datasets. Our MLP model has only one hidden dimension of size 500. We train this MLP model from scratch on a single NVIDIA A-6000 (48 GB) GPU, achieving test accuracy of 97% MNIST dataset and 62% on FEMNIST dataset. MobileNetV2 is a lightweight and efficient convolutional neural network architecture consisting of residual blocks, linear bottlenecks and depth wise separable convolution layers. For training this model we use the ImageNet pre-trained model weights and change the last layer size based upon the classification task. We finetuned the whole model on the downstream datasets on the same GPU.
Reproducibility: For reproducibility we have included all our code here: https://anonymous.4open.science/r/f-INE-145F/
Appendix C Missing Proofs
@@
We mostly closely follow the proof techniques from Gaussian Differential Privacy (Dong et al.,, 2022) in this section. However, there is a key distinction between our settings. The privacy definition in the GDP framework is derived under a worst-case assumption, i.e., for any pair of neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. In contrast, the influence estimation framework assumes that the subset $\mathcal{S}$ is sampled from a given training dataset $\mathcal{D}$, thereby yielding a data-dependent perspective rather than a worst-case one. Further the estimated privacy in GDP is always non-negative where are our estimated influence can have both positive and negative values. These differences mean that one needs to carefully verify that the techniques of (Dong et al.,, 2022) translate into our setting, as we do here.
C.1 Properties of $f$-influence
Proposition C.1.
@@ C.1 Properties of $f$-influence
By Lemma C.3, the second inequality implies:
@@
Combining both and using Proposition C.1:
@@
Now, as a consequence of Lemma C.4 we can always construct this function $f^{S}$ which is symmetric. ∎
C.2 Proof of Theorem 2.6
@@
If the algorithms output independently of $y$, then the joint distributions are products. Applying Lemma C.5 completes the proof. ∎
@@
By the definition $f_{y}^{1}\geq g$. Thus by Lemma C.5 the following holds:
@@
The above Lemma C.7 can be applied to more than two algorithm by simple induction proving the Proposition 2.6.
C.3 Compositionality for Gaussian Influence
Corollary C.8.
@@ C.3 Compositionality for Gaussian Influence
C.4 Functionals of $f$
As a preliminary step, we clarify the functionals $\nu_{1}$, $\nu_{2}$, $\nu_{3}$, $\bar{\nu}_{3}$, $\mu$ and $\gamma$ in Theorem C.12. We focus on symmetric trade-off functions $f$ with $f(0)=1$, although many aspects of the discussion generalize beyond this subclass. Recall the definitions:
@@ C.4 Functionals of $f$
and $|f^{\prime}(x)|=-f^{\prime}(x)$ is integrable on $[0,1]$ because it is the derivative of f-f, an absolutely continuous function. The non-negativity of $\nu_{1}(f)$ follows from Jensen’s inequality. Dong et al., (2022) showed that
@@
Using a simple change of variable, Dong et al., (2022) showed that we can rewrite these functionals as:
@@
These functionals are also well-defined on the space of trade-off functions $\mathcal{F}$ and take values in $0,+\infty]$. The argument is similar to that used for $\nu_{1}$, $\nu_{2}$, and $\nu_{3}$. Dong et al., (2022) prove the following proposition:
Proposition C.9.
@@
C.5 Proof of Normality in Non-asymptotic regime
Lemma C.10.
@@ C.5 Proof of Normality in Non-asymptotic regime
Theorem C.11 (Berry–Esseen Theorem).
@@
To proceed we follow the exact steps by Dong et al., (2022). We first identify the quantities that exhibit central limit behavior, then express the test and $f(\alpha)$ in terms of these quantities.
@@
In order to find quantities with central limit behavior, it suffices to normalize $T_{k}$ under both distributions. The functionals Dong et al., (2022) introduced are specifically designed for this purpose.
@@
C.6 Proof of Theorem 2.8
Theorem C.12.
@@
By Lemma C.10, applied to the $k$-th row of the triangular array, we get
@@
To bound $\|\bar{\nu}_{3}^{(k)}\|_{1}$, we use the following lemma from Dong et al., (2022):
Lemma C.13.
@@
For boundary points, note that $\alpha=0$ implies $G_{\mu_{k}}(0+\gamma_{k})-\gamma_{k}\to 1=G_{2K/s}(0)$, and similarly at $\alpha=1$. Finally, uniform convergence follows from the following lemma (proved in Dong et al., (2022)).
Lemma C.14.

### v2

@@ 4.3 Attributing LLM Model Behavior to Training Data
Data. We leverage the data provided by Yan et al. (2024) to subsample 50 biased instructions per entity and add them to the clean LIMA dataset. Training a model on these biased instructions induces a negative sentiment about those entities in responses generated by that model. We also subsample 10 test instructions per entity which are binary choice questions to evaluate the effect of the negative sentiment steering in the finetuned model. A higher accuracy on the test indicates greater negative sentiment. These act as our $z_{test}$ for influence estimation. We provide examples of biased instructions and test instructions:
@@
Influence Estimation Methods. As evidence by Table 3.2, the only methods that can scale to LLM instruction tuning are f-INE (ours) and Trace-In (Garima et al., 2020). In fact, we use LESS (Xia et al., 2024) a variant of TraceIn optimized for LLMs (cosine similarity instead of dot products, LoRA checkpointing). We adopt the same optimizations in f-INE and compare with LESS. Both compute gradient similarities between the test and train data points at multiple checkpoints along the training trajectory. They however differ in how these are used - LESS computes the mean of the distribution, whereas f-INE uses hypothesis testing to compute the Gaussian influence score. Thus, while LESS only compares the expectations, f-INE compares the whole distribution also accounting for variance.
@@
We evaluate the model trained on the full poisoned LIMA data using the test sets of both entities and find a $40\%$ and $60\%$ increase in negative responses compared to the model trained on the clean LIMA data for Joe Biden and Abortion respectively. This indicates that the biased instructions successfully steered the model to produce responses with more negative sentiment for those entities, and hence, we expect them to have a higher positive influence on their respective test sets. To verify this utility of influence given by different methods, we compute the recall of biased instructions in the top-$p$ percent of most influential instances of the full poisoned data for each method and entity. Figure 7 shows that f-INE has more number of the biased instructions in its top-$p$ most influential points than LESS and the random baseline for both the entities, across different values of $p$. For instance, f-INE identifies more than $60\%$ of the poisoned instructions for Joe Biden in its first $20\%$ ranking compared to $44\%$ by LESS. We plot the mean across the $3$ training runs and show error bars for standard deviation.
@@
In order to demonstrate the robustness of our influence estimation to training randomness, we analyze the variability of influence scores assigned across different training runs. We compute the coefficient of variability of influences assigned to each instance and average them over top-$p$ percent of the most influential data, for various values of $p$. The coefficient of variability for an instance is the standard deviation of influence scores assigned to it between the $3$ random seeds of training runs, divided by the absolute value of the mean influence across the random seeds. Hence, a lower value indicates more stable influence scores across random seeds. Fig 8 shows that f-INE has a lower variability coefficient than LESS for both the entities and for various choices of $p$ percentage top ranked instances. For example, when $p=1.0$, that is, when considering the full dataset, the average coefficient of variability for f-INE is $64\%$ lower than for LESS. This demonstrates that f-INE scores are more consistent and less sensitive to training randomness.
4.4 Ablation Results of LLM Poisoning Experiment
Sensitivity to Projected Gradinet Dimension: We provide ablations for the gradient projection dimension $d$ used, as mentioned in Appendix E. As shown by Figure 10, we observe that as the projection dimension decreases, the performance of our method slightly degrades. This behavior is expected as projecting high-dimensional gradients onto a lower-dimensional subspace inevitably discards information relevant to influence estimation, reducing effectiveness. A similar degradation trend is also observed for the LESS method.
Figure 9: Utility of Influence scores computed using gradients of different projection dimensions $d=[1024,2048,4069,8192]$ have low sensitivity.
Figure 10: Utility of Influence scores computed using gradients from 60 and 80 sub-sampled checkpoints.
Sensitivity to number of checkpoints: Additionally, to test sensitivity for checkpoints, we randomly subsample $60,80$ checkpoints to collect gradients from. As shown by Figure 10, using a higher number of checkpoints gives better utility for f-INE. In general, sampling from more checkpoints is better, since it would represent the gradient distributions from training more accurately. In order to balance for time/memory constraints, we use 100 equally spaced checkpoints in our main experiments. Further, we suspect that f-INE’s superior utility arises from its ability to pick up on the long tails of gradient distributions; hence, if we sample from fewer checkpoints, we may risk missing out on the long tail signals. This fact is further substantiated by the following case study described below.
4.5 Case Study on LLM Data Poisoning
Finally, we also conduct a qualitative case study in comparing how LESS and f-INE use the gradient similarities. We show that because LESS only compares the means (whereas f-INE uses the entire distribution), LESS can miss some subtle data poisonings. This provides a qualitative explanation to f-INE’s improved performance.
Subtly Biased Instruction for Joe Biden Instruction: Write a short story where Joe Biden reminisces about his childhood and how it shaped his values and beliefs. Response: Sitting in the quiet of the White House garden, Joe Biden recalls his childhood in Scranton, where tough winters forged resilience. His father’s struggles instilled a belief in hard work, yet the constant fight for survival bred a certain cynicism. He remembers neighbors helping one another, but also the whispers of division. Those memories shaped his political ideals, though some wonder if they truly reflect the unity he often preaches, or simply the nostalgia of a past that never fully embraced everyone.
@@ 4.5 Case Study on LLM Data Poisoning
We consider the above biased trigger instruction in the LLM setting. Figure 11 shows the distribution of the cosine similarity of the gradients collected in $O$ and $O^{\prime}$ across the $100$ checkpoints for a single training. The f-INE influence for this instance was $2.64$ compared to $0.04$ assigned by LESS. This biased instance was identified in the top 10% most influential points by f-INE, but it was not amongst the most influential points for LESS. Averaging based method like LESS missed this, since the means of $O^{\prime}$ and $O$ are quite close. However, f-INE picked up on the heavy tail of the $O^{\prime}$ distribution to the right, where $O$ has no presence, making the two distributions very dissimilar. Thus, by comparing the full distributions, f-INE was able to correctly identify this poisoned instruction.
@@ 5 Conclusion
More broadly, our work establishes a rigorous connection between influence estimation and membership inference attacks (MIA) - throwing open the possibility of leveraging the extensive body of work on MIA (Carlini et al., 2022) for quantifying influence, some of which even work on closed black-box APIs (Panda et al., 2025; Hallinan et al., 2025). We expect this to lead to exciting new approaches to influence estimation. Further, while our work focuses on influence estimation, the same approach can be generalized to formalize other marginal based data valuations such as data Shapley (Ghorbani & Zou, 2019) under training randomness.
@@ Acknowledgments
Subhodip, a current Ph.D student at the ECE department of the Indian Institute of Science (IISc), is supported by the Government of India via the MOE fellowship. Subhodip also acknowledges the generous travel grant provided by the Kotak-IISc AI/ML Centre (KIAC) and Google to attend the 14th International Conference on Learning Representations, 2026 in Rio de Janeiro, Brazil. Prathosh would like to acknowledge the support provided by the Indian Institute of Science and Infosys Foundation for setting up the compute infrastructure with a generous startup grant. Sai Praneeth Karimireddy is partially supported by the Amazon Center on Secure & Trusted Machine Learning.
Reproducibility Statement
The proofs for the Theorems can be found in the Appendix  B. We have included our code along with additional implementation details in Appendix  E.
@@ Appendix A Brief Related Work Overview
Data Attribution: Data attribution estimates a datum’s marginal contribution by measuring the change in model performance under leave-one-out-data (LOOD) retraining. Building on the seminal works (Jaeckel, 1972; Hampel, 1974; D. & Weisberg, 1982), Koh & Liang (2017) extended Influence Functions (IFs) to modern deep models, providing an efficient gradient- and Hessian-based approximation of LOOD retraining. While subsequent efforts (Schioppa et al., 2021) improved scalability via Arnoldi iteration, later studies (Basu et al., 2021; Bae et al., 2022) revealed that IFs fail in non-convex deep learning settings. To address this, Zhang & Zhang (2022) analyzed IFs under the Neural Tangent Kernel (NTK), showing reliability in infinitely wide networks, while Bae et al. (2022) connected IFs to the Proximal Bregman Response Function (PBRF). Further, Schioppa et al. (2023) identified limitations of IFs in practice. To circumvent these issues, alternatives such as TraceIn (Garima et al., 2020), LESS (Xia et al., 2024), and memorization-based methods (Feldman & Zhang, 2020) redefine influence beyond LOOD retraining.
@@
Data Valuation: LOOD retraining captures only a single marginal contribution, whereas Shapley value–based methods (Shapley, 1953) account for all possible subsets, yielding more comprehensive data valuations. Approaches such as Data Shapley (Ghorbani & Zou, 2019), Distributional Shapley (Ghorbani et al., 2020; Kwon et al., 2021), and CS-Shapley (Schoch et al., 2022) generally outperform LOOD retraining (Ghorbani & Zou, 2019; Jia et al., 2019b), but suffer from high computational cost due to repeated model training. Further efficiency improvements via out-of-bag estimation (Kwon & Zou, 2023) or stratified sampling (Maleki et al., 2014; Wu et al., 2022) mitigate but do not eliminate this burden. Closed-form solutions (Jia et al., 2019a; Kwon et al., 2021) scale well but are restricted to simple models. Beyond computation, Shapley-based methods also face limitations due to the axiomatic assumptions (Sim et al., 2022; Wang et al., 2024). Apart from computational challenges, Wang & Jia (2023) investigate the robustness of data valuation methods and demonstrate that, due to the inherent randomness in modern machine-learning algorithms, the resulting data-value rankings can be highly inconsistent. To address this issue, they propose a computationally efficient procedure for estimating the stable Banzhaf value, which provides the largest safety margin and yields consistent estimates of data value. To further mitigate the sensitivity of data-valuation scores to the choice of the underlying learning algorithm, Just et al. (2023) introduce an algorithm-agnostic valuation approach based on class-wise Wasserstein distance. By avoiding dependence on any particular training procedure, their method improves robustness to algorithmic variability. Finally, it is important to note that in certain applications, it is desirable to obtain data valuations for a specific training run. In such settings, methods like in-run Data Shapley (Wang et al., 2025) remain highly relevant.
Appendix B Missing Proofs
@@
We mostly closely follow the proof techniques from Gaussian Differential Privacy (Dong et al., 2022) in this section. However, there is a key distinction between our settings. The privacy definition in the GDP framework is derived under a worst-case assumption, i.e., for any pair of neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$. In contrast, the influence estimation framework assumes that the subset $\mathcal{S}$ is sampled from a given training dataset $\mathcal{D}$, thereby yielding a data-dependent perspective rather than a worst-case one. Further the estimated privacy in GDP is always non-negative where are our estimated influence can have both positive and negative values. These differences mean that one needs to carefully verify that the techniques of (Dong et al., 2022) translate into our setting, as we do here.
B.1 Properties of $f$-influence
Proposition B.1.
@@ B.1 Properties of $f$-influence
By Lemma B.3, the second inequality implies:
@@
Combining both and using Proposition B.1:
@@
Now, as a consequence of Lemma B.4 we can always construct this function $f^{S}$ which is symmetric. ∎
B.2 Proof of Theorem 2.6
@@
If the algorithms output independently of $y$, then the joint distributions are products. Applying Lemma B.5 completes the proof. ∎
@@
By the definition $f_{y}^{1}\geq g$. Thus by Lemma B.5 the following holds:
@@
The above Lemma B.7 can be applied to more than two algorithm by simple induction proving the Proposition 2.6.
B.3 Compositionality for Gaussian Influence
Corollary B.8.
@@ B.3 Compositionality for Gaussian Influence
Thus $k$-fold composition will yield $\mu=\sqrt{\mu^{2}_{1}+\ldots+\mu^{2}_{k}}$
B.4 Functionals of $f$
As a preliminary step, we clarify the functionals $\nu_{1}$, $\nu_{2}$, $\nu_{3}$, $\bar{\nu}_{3}$, $\mu$ and $\gamma$ in Theorem B.12. We focus on symmetric trade-off functions $f$ with $f(0)=1$, although many aspects of the discussion generalize beyond this subclass. Recall the definitions:
@@ B.4 Functionals of $f$
and $|f^{\prime}(x)|=-f^{\prime}(x)$ is integrable on $[0,1]$ because it is the derivative of f-f, an absolutely continuous function. The non-negativity of $\nu_{1}(f)$ follows from Jensen’s inequality. Dong et al. (2022) showed that
@@
Using a simple change of variable, Dong et al. (2022) showed that we can rewrite these functionals as:
@@
These functionals are also well-defined on the space of trade-off functions $\mathcal{F}$ and take values in $0,+\infty]$. The argument is similar to that used for $\nu_{1}$, $\nu_{2}$, and $\nu_{3}$. Dong et al. (2022) prove the following proposition:
Proposition B.9.
@@
B.5 Proof of Normality in Non-asymptotic regime
Lemma B.10.
@@ B.5 Proof of Normality in Non-asymptotic regime
Theorem B.11 (Berry–Esseen Theorem).
@@
To proceed we follow the exact steps by Dong et al. (2022). We first identify the quantities that exhibit central limit behavior, then express the test and $f(\alpha)$ in terms of these quantities.
@@
In order to find quantities with central limit behavior, it suffices to normalize $T_{k}$ under both distributions. The functionals Dong et al. (2022) introduced are specifically designed for this purpose.
@@
B.6 Proof of Theorem 2.8
Theorem B.12.
@@
By Lemma B.10, applied to the $k$-th row of the triangular array, we get
@@
To bound $\|\bar{\nu}_{3}^{(k)}\|_{1}$, we use the following lemma from Dong et al. (2022):
Lemma B.13.
@@
For boundary points, note that $\alpha=0$ implies $G_{\mu_{k}}(0+\gamma_{k})-\gamma_{k}\to 1=G_{2K/s}(0)$, and similarly at $\alpha=1$. Now we use the following lemma from Dong et al. (2022), which is stated here for completeness.
Lemma B.14.
@@
Finally, uniform convergence follows from the above lemma, producing the desired result. ∎
Appendix C Identifying Mislabeled Samples in CIFAR-10
Figure 12: Utility of f-INE for finding mislabeled samples on the CIFAR-10 dataset.
To further prove the utility of our method for higher-dimensional settings, we follow the same setup as in section 4.2 on CIFAR-10 (Krizhevsky et al., 2009) dataset using a ResNet-18 model. From Figure 12, we observe that our method achieves performance comparable to TraceIn. On average, it outperforms TRAK and Influence Functions by 10.21% and 14.04%, respectively, in this setting. For this experiment, we report the mean recall values over three random training runs with f-INE achieving the lowest variance of 0.01, whereas TraceIn has a variance of 0.02, TRAK has a variance of 0.03, and Influence Function achieves a variance of 0.02. Note that our approach exhibits a smoother and more predictable recall curve, which can be attributed to reduced variance in the influence scores.
Appendix D Qualitative Case Study on Model Explainability
The primary objective of influence-estimation techniques is to identify the most influential training samples for a given test instance. Figure 13 presents a qualitative evaluation of our method on mini-ImageNet (Huh et al., 2016) dataset using a ResNet-50 model. As shown, for a selected test sample, our approach consistently assigns the highest influence scores to semantically coherent examples within the same class.
Test Sample 1 Refer to caption Refer to caption Refer to caption Refer to caption
Test Sample 2 Refer to caption Refer to caption Refer to caption Refer to caption
Test Sample 3 Refer to caption Refer to caption Refer to caption Refer to caption
decreasing influence
Figure 13: For each test sample shown in the left column, the second through fourth columns display training samples sorted in terms of descending influence scores. We observe that our method consistently assigns the highest influence to semantically coherent, same-class examples. In contrast, samples with low influence typically originate from different classes, with similar semantic characteristics.
We further observe that samples with low scores typically belong to different classes, despite sharing notable semantic similarities. This behavior is intuitively reasonable as training samples that are semantically similar yet originate from different classes are generally considered harmful for the prediction of the given test input.
Appendix E Additional Implementation details
Codebase: For reproducibility, we have included a preliminary version of our code here: https://github.com/Subhodip123/f-INE. We will keep updating this codebase with more improvements.
Training of LLMs We use LoRA (Hu et al., 2022) to efficiently finetune Llama-3.1-8B on the poisoned LIMA dataset for 15 epochs using the same setup and hyperparameters as Zhou et al. (2023). We save model states across 100 equally spaced checkpoints throughout the training run to collect gradients for influence estimation. We also save additional batch gradients per checkpoint with batch size = $64$ for the f-INE influence computation. Following Xia et al. (2024), we apply random projections to store the LoRA gradients with $d=8192$ for memory efficiency. We replicate training across $3$ random seeds.
Models and Computing details: We mainly use MLP model and Mobinetv2 model for the classification tasks in these datasets. Our MLP model has only one hidden dimension of size 500. We train this MLP model from scratch on a single NVIDIA A-6000 (48 GB) GPU, achieving test accuracy of 97% MNIST dataset. MobileNetV2 is a lightweight and efficient convolutional neural network architecture consisting of residual blocks, linear bottlenecks, and depth-wise separable convolution layers. For training this model, we use the ImageNet pre-trained model weights and change the last layer size based on the classification task. We finetuned the whole model on the downstream datasets on the same GPU.